OpenStudy (anonymous):

The minimum and maximum temperature on a cold day in Lollypop Town can be modeled by the equation below: 2|x − 6| + 14 = 38 What are the minimum and maximum temperatures for this day? x = −9, x = 21 x = −6, x = 18 x = 6, x = −18 No solution I got B, is that right ?

OpenStudy (anonymous):

2|x − 6| + 14 = 38 2|x − 6| = 24 |x − 6| = 12 x-6 = 12 OR x-6 = -12 x = 18 or x = -6 Yes, I get B too

OpenStudy (anonymous):

Thank you!